Find the p-value for the t-test of Ho: µ = 4 vs. Ha: µ > 4 where n = 11 and t = 2.646. Report 4 decimal places.
Answer #1

Solution :

Given that,

This is the right tailed test .

t = 2.646

n = 11

df = n - 1 = 11 - 1 = 10

p-value = 0.0122

P-value for the test is = 0.0122